# MGSR: 2D/3D Mutual-boosted Gaussian Splatting for High-fidelity Surface Reconstruction under Various Light Conditions

### v1

@@ 3.3 Surface reconstruction with 2D-GS
$Z=\frac{\sum_{i}\omega_{i}z_{i}}{\sum_{i}\omega_{i}+\epsilon},$ (7)
@@
$T_{i}=\prod_{j=1}^{i-1}\left(1-\alpha_{j}\right),$ (8)
@@
$N(x,y)=\frac{\nabla_{x}Z\times\nabla_{y}Z}{|\nabla_{x}Z\times\nabla_{y}Z|}.$ (10)
Considering that our task is to reconstruct the surface meshes under various light conditions, a weighted normal loss is employed, which encourages the model to primarily learn the normals of the surfaces rather than those of the entire scenes. The normal loss from the camera’s perspective will be determined based on the division of the mask into foreground and background, with a coefficient $\gamma$ applied to weight the contributions. Two TV losses $\mathcal{L}_{\text{d-TV}}$ and $\mathcal{L}_{\text{n-TV}}$ are introduced as smooth terms on both rendered depths and normals. The overall loss of the 2D-GS branch consists of a weighted combination:
@@
where $\lambda_{3}$, $\lambda_{4}$, and $\lambda_{5}$ denote to the weights, and $\mathcal{L}_{\text{render}}$ is defined as Eq. 5.
Yet, the 2D-GS branch has only accomplished modeling of illuminated scenes and relatively reliable depth estimation without eliminating the influence of light conditions, which is inconsistent with real unlit scenes. To address this issue, an alternating optimization approach is devised after the warm-up stage of both the 2D-GS and 3D-GS branches.
@@ 3.4 Alternating optimization of 2D & 3D Gaussians
Figure 3: Geometry enhancement in 3D-GS branch for realistic rendering through our mutual-boosted optimization.
Given that the warm-up 3D-GS branch has performed a preliminary illumination decomposition in Sec. 3.2, and the warm-up 2D-GS branch has provided an initial estimate of the depth maps for the scenes under various light conditions in Sec. 3.3, our alternating optimization approach utilizes the transmitted colors from the 3D-GS branch to supervise the 2D-GS branch for surface reconstruction, while the depth maps from the 2D-GS branch provides geometry information for the 3D-GS branch for better illumination decomposition. Specifically, since the illumination decomposition in the 3D-GS branch is unsupervised, the transmitted components from 3D-GS branch might not be completely disentangled from the reflected components after the warm-up stage, a weighted rendering loss is introduced for rendering supervision for the 2D-GS branch.
Our goal is for the images rendered by the 2D-GS branch to not only approximate the transmitted images but also to retain a slight similarity to the illuminated images to enhance the stability of alternating optimization, as calculated by:
@@
where $\mathcal{L}_{\text{render-m}}$ represents the mutual rendering loss in alternating optimization, and $\lambda_{6}$ is the weight. $\mathcal{L}_{\text{mutual}}$ denotes the rendering loss between the 2D-GS branch rendered images and the transmitted images from the 3D-GS branch, $\mathcal{L}_{\text{2D-render}}$ is the rendering loss between the 2D-GS branch rendered images and GT images, and both of the rendering loss are computed by Eq. 5. Moreover, both TV losses on depths ($\mathcal{L}_{\text{d-TV-m}}$) and normals ($\mathcal{L}_{\text{n-TV-m}}$) in 2D-GS branch are retained, but GT images are replaced with the transmitted images. In the alternating optimization stage, the loss function of 2D-GS branch will be promoted to:
@@
where $\gamma$ is the coefficient that balances the contribution of the foreground and background. $\lambda_{2}$, $\lambda_{3}$, $\lambda_{4}$ are same as the warm-up stage.
Scene geometry is essential for realistic physically-based rendering, and geometry estimation on Gaussian primitives is difficult due to the discrete structures. Therefore, on the other hand, a depth loss between 2D-GS and 3D-GS branches is introduced to improve the geometry estimation for the illumination decomposition of 3D-GS branch (as shown in Fig. 3). The calculation of depth maps $Z_{\text{3D}}$ in 3D-GS branch is the same as Eq. 7. To encourage the alternating optimization to focus on the foreground part of the input images, masks are used to avoid the calculation of background by with a weight coefficient $\gamma$. The depth loss $\mathcal{L}_{Z}$ is computed by:
@@
where $\mathcal{L}_{2}$ denotes to the L2 loss, and the $Z_{\text{2D}}$ is estimated depth maps $Z$ in the 2D-GS branch.
@@
where the rendering loss of the 3D module $\mathcal{L}_{\text{3D}}$ is calculated by Eq. 3 (rendering loss only), the normal loss of the 2D module $\mathcal{L}_{\text{n}}$ is computed by Eq. 9, Eq. 12, $w_{\text{2D}}$, $w_{\text{3D}}$, and $w_{\text{depth-mutual}}$ are the weights.
@@
Figure 4: Visual comparisons on the OmniObject3D dataset [27].
Figure 5: Visual comparisons on the Shiny Blender dataset [23].
@@ 4.1 Datasets and Evaluation Metrics
4.2 Results
Object-level. Three objects from the Shiny Blender with reflections and 30 objects from OmniObject3D with highlights are conducted on all methods for comparison. Table 1 and Table 2 present quantitative results on both NVS and SR. MGSR surpasses all baselines in terms of the PSNR metric for NVS and the NC metric for SR. As shown in visualization comparisons of Figure 4 and Figure 5, although NeuS2 exhibits superior CD, its smoothness on surfaces with reflections is significantly poor. To address this issue, NC is introduced as an evaluation metric for reconstruction, overcoming the limitations of CDs, which fail to capture surface holes or bumps. 2D-GS produces excessive faces on reconstructed meshes on the exterior of objects while failing to accurately reconstruct thin surfaces. GOF tends to reconstruct rough or incomplete surfaces when handling lit surfaces. MGSR visually outperforms all baselines, resulting in the best NC, with smooth surfaces and accurate color modeling. Table 3 and Figure 6 show the quantitative and visual results compared to GS-based SR methods on DTU dataset. Under varying light conditions, MGSR successfully reconstructs realistic and intact surfaces compared to 2D-GS and GOF.
Scene-level. As shown in Figure 7, two in-the-wild scenes from Ref-NeRF [23] and TnT [29] are used for visual comparisons. More comparisons are shown in the Appendix. Previous GS-based methods fail to effectively reconstruct glass or mirror surfaces, resulting in damaged and inaccurate surfaces. In contrast, MGSR successfully reconstructs meshes in mirror material, driven by our devised 3D-GS branch and mutual-boosted optimization, which perform illumination decomposition and guide the 2D-GS branch for enhanced SR.
@@ 4.2 Results
Optimization Time. In Table 1 and Table 2, we additionally present the optimization time for all compared methods. By eliminating the dependence on BRDF, MGSR is faster than all illumination decomposition baselines. Among methods supporting both NVS and SR, 2D-GS is the fastest. MGSR, supported by an auto-stop warm-up strategy, outperforms GOF and achieves a comparable speed with 2D-GS. Specifically, each of the two branches in MGSR has an average warm-up optimization time of around 3.5-4 minutes. The mutual-boosted optimization time is approximately 8-9 minutes.
@@ 5 Ablation studies
All ablation studies (Table 4) are conducted on the OmniObject3D dataset due to its various light conditions, with the experimental setup consistent with Section 4. More additional results are provided in the Appendix.
Auto-stop Warm-up. Experimental results (Models G-J) show that the auto-stop strategy during warm-up outperforms both MVS and SR, demonstrating the effectiveness of the auto-stop strategy in preventing overfitting during the warm-up stage. Bidirectional Back-propagation (BP). In alternating optimization (Section 3.4), the outputs from one branch are utilized to supervise the other branch. During the supervision, the gradient of loss can be back-propagated to one branch (unidirectional BP) or both branches (bidirectional BP). Experimental results (Models K and M) show that unidirectional BP outperforms bidirectional BP in terms of both NVS and SR results. It is mainly due to the bidirectional BP influences optimization of the branch which provides supervision. Mutual-boosted Iterations. It is found that when the mutual-boosted iterations exceed 20k, there is no significant improvement in the NVS and SR metrics (Models L and M). Therefore, 20k iterations are optimal for MGSR. Weights of Loss. The cases without mutual-boosted optimization are firstly validated, where only the 3D-GS branch is used for NVS or the 2D-GS branch for SR (Models A and B). The results were inferior compared to the case with mutual-boosted optimization. Furthermore, it is found that the branch with a higher weight (Models C and D) during the mutual-boosted optimization stage tends to favor the task it excels at, leading to better performance. To balance both NVS and SR, the weights of the two branches should be equal.
@@ 6 Conclusion
In this study, we introduce MGSR, a 2D/3D Mutual-boosted Gaussian splatting for SR that enhances both rendering quality and 3D reconstruction accuracy under various light conditions. Moreover, the auto-stop strategy is proposed in the warm-up stage, while geometry-guided illumination decomposition is devised in the mutual-boosted optimization stage. Extensive experiments across various synthetic and real-world datasets at both object and scene levels have validated the superiority of MGSR.
Limitations and Future Work. Although MGSR may achieve reliable reconstruction under strong light or reflective conditions, overly smooth surfaces may occur in low-light scenarios, leading to the absence of details. We will address this issue by incorporating exposure compensation for the input images.

### v2

@@ 3.3 Surface reconstruction with 2DGS
$\small Z=\frac{\sum_{i}\omega_{i}z_{i}}{\sum_{i}\omega_{i}+\epsilon},$ (7)
@@
$\small T_{i}=\prod_{j=1}^{i-1}\left(1-\alpha_{j}\right),$ (8)
@@
$N(x,y)=\frac{\nabla_{x}Z\times\nabla_{y}Z}{\|\nabla_{x}Z\times\nabla_{y}Z\|}.$ (10)
Considering that our task is to reconstruct the surface meshes under various light conditions, a weighted normal loss is employed, which encourages the model to primarily learn the normals of the surfaces rather than those of the entire scenes. The normal loss from the camera’s perspective will be determined based on the division of the mask into foreground and background, with a coefficient $\gamma$ applied to weight the contributions. Two TV losses $\mathcal{L}_{\text{d-TV}}$ and $\mathcal{L}_{\text{n-TV}}$ are introduced as smooth terms on both rendered depths and normals. The overall loss of the 2DGS branch consists of a weighted combination:
@@
where $\lambda_{3}$, $\lambda_{4}$, and $\lambda_{5}$ represent the weights, and $\mathcal{L}_{\text{render}}$ is defined in Equation 5.
Yet, the 2DGS branch has only accomplished modeling of illuminated scenes and relatively reliable depth estimation without eliminating the influence of light conditions, which is inconsistent with real unlit scenes. To address this issue, an alternating optimization approach is devised after the warm-up stage of both the 2DGS and 3DGS branches.
@@ 3.4 Alternating optimization of 2D & 3D Gaussians
Figure 3: Geometry enhancement in 3DGS branch for realistic rendering through our mutual-boosted optimization.
Given that the warm-up 3DGS branch has performed a preliminary illumination decomposition in Section 3.2, and the warm-up 2DGS branch has provided an initial estimate of the depth maps for the scenes under various light conditions in Section 3.3, our alternating optimization approach utilizes the transmitted colors from the 3DGS branch to supervise the 2DGS branch for surface reconstruction, while the depth maps from the 2DGS branch provides geometry information for the 3DGS branch for better illumination decomposition.
Specifically, since the illumination decomposition in the 3DGS branch is unsupervised, the transmitted components from 3DGS branch might not be completely disentangled from the reflected components after the warm-up stage, a weighted rendering loss is introduced for rendering supervision for the 2DGS branch. Our goal is for the images rendered by the 2DGS branch to not only approximate the transmitted images but also to retain a slight similarity to the illuminated images to enhance the stability of alternating optimization, as calculated by:
@@
where $\mathcal{L}_{\text{render-m}}$ represents the mutual rendering loss in alternating optimization, and $\lambda_{6}$ is the weight. $\mathcal{L}_{\text{mutual}}$ denotes the rendering loss between the 2DGS branch rendered images and the transmitted images from the 3DGS branch, $\mathcal{L}_{\text{2D-render}}$ is the rendering loss between the 2DGS branch rendered images and the GT images, and both of the rendering loss are computed by Equation 5. Moreover, both TV losses on depths ($\mathcal{L}_{\text{d-TV-m}}$) and normals ($\mathcal{L}_{\text{n-TV-m}}$) in 2DGS branch are retained, but GT images are replaced with the transmitted images. In the alternating optimization stage, the loss function of 2DGS branch will be promoted to:
@@
where $\gamma$ is the coefficient that balances the contribution of the foreground and background. The weights $\lambda_{2}$, $\lambda_{3}$, and $\lambda_{4}$ are the same as the warm-up stage.
Figure 4: Visual comparisons on the OmniObject3D dataset [22].
Figure 5: Visual comparisons on the Shiny Blender dataset [21].
Scene geometry is essential for realistic physically-based rendering, and geometry estimation on Gaussian primitives is difficult due to the discrete structures. Therefore, on the other hand, a depth loss between 2DGS and 3DGS branches is introduced to improve the geometry estimation for the illumination decomposition of 3DGS branch (as shown in Figure 3). The calculation of depth maps $Z_{\text{3D}}$ in 3DGS branch is the same as Equation 7. To encourage the alternating optimization to focus on the foreground part of the input images, masks are used to avoid the calculation of background by with a weight coefficient $\gamma$. The depth loss $\mathcal{L}_{Z}$ is computed by:
@@
where $\mathcal{L}_{2}$ denotes the L2 loss, and $Z_{\text{2D}}$ is the estimated depth map $Z$ in the 2DGS branch.
@@
where the losses of the 3D module $\mathcal{L}_{\text{3D}}$ and the 2D module $\mathcal{L}_{\text{2D}}$ are calculated by Equation 6 and Equation 13, respectively, and $w_{\text{2D}}$, $w_{\text{3D}}$, and $w_{\text{depth-mutual}}$ are the weights. See Section 5 for their settings.
@@ 4.1 Datasets and evaluation metrics
4.2 Results
Object-level data. Three objects from the Shiny Blender with reflections and 30 objects from OmniObject3D with highlights are conducted on all methods for comparison. Table 1 and Table 2 present quantitative results on both NVS and SR. MGSR surpasses all baselines in terms of the PSNR metric for NVS and the NC metric for SR. As shown in visualization comparisons of Figure 4 and Figure 5, although NeuS2 exhibits superior CDs, its smoothness on surfaces with reflections is significantly poor. To address this issue, NC is introduced as an evaluation metric for reconstruction, overcoming the limitations of CDs, which fail to capture surface holes or bumps. 2DGS produces excessive faces on reconstructed meshes on the exterior of objects while failing to accurately reconstruct thin surfaces. GOF tends to reconstruct rough or incomplete surfaces when handling lit surfaces. MGSR visually outperforms all baselines, resulting in the best NC, with smooth surfaces and accurate color modeling. Table 3 and Figure 6 show the quantitative and visual results compared to GS-based SR methods on DTU dataset. Under varying light conditions, MGSR successfully reconstructs realistic and intact surfaces compared to 2DGS and GOF.
Scene-level data. As shown in Figure 7, two in-the-wild scenes from Ref-NeRF [21] and TnT [10] are used for visual comparisons. More comparisons are shown in the Appendix. Previous GS-based methods fail to effectively reconstruct glass or mirror surfaces, resulting in damaged and inaccurate surfaces. In contrast, MGSR successfully reconstructs meshes in mirror material, driven by our devised 3DGS branch and mutual-boosted optimization, which perform illumination decomposition and guide the 2DGS branch for enhanced SR.
Optimization time. In Table 1 and Table 2, we additionally present the optimization time for all compared methods. By eliminating the dependence on BRDF, MGSR is faster than all illumination decomposition baselines. Among methods supporting both NVS and SR, 2DGS is the fastest. MGSR, supported by an auto-stop warm-up strategy, outperforms GOF and achieves a comparable speed with 2DGS. Specifically, each of the two branches in MGSR has an average warm-up optimization time of around 3.5-4 minutes. The mutual-boosted optimization time is approximately 8-9 minutes.
@@ 5 Ablation studies
All ablation studies (Table 4) are conducted on the OmniObject3D dataset due to its various light conditions, with the experimental setup consistent with Section 4. Additional results are provided in the Appendix.
Auto-stop warm-up. Experimental results (Models G-J) show that the auto-stop strategy during warm-up outperforms both MVS and SR, demonstrating the effectiveness of the auto-stop strategy in preventing overfitting during the warm-up stage. Bidirectional back-propagation (BP). In alternating optimization (Section 3.4), the outputs from one branch are utilized to supervise the other branch. During the supervision, the gradient of loss can be back-propagated to one branch (unidirectional BP) or both branches (bidirectional BP). Experimental results (Models K and M) show that unidirectional BP outperforms bidirectional BP in terms of both NVS and SR results. It is mainly due to the bidirectional BP influences optimization of the branch which provides supervision. Mutual-boosted iterations. We observed that when the mutual-boosted iterations exceed 20k, there is no significant improvement in the NVS and SR metrics (Models L and M). Therefore, 20k iterations are optimal for MGSR. Loss weights. The cases without mutual-boosted optimization are firstly validated, where only the 3DGS branch is used for NVS or the 2DGS branch for SR (Models A and B). The results were inferior compared to the case with mutual-boosted optimization. Furthermore, it is found that the branch with a higher weight (Models C and D) during the mutual-boosted optimization stage tends to favor the task it excels at, leading to better performance. To balance both NVS and SR, the weights of the two branches should be equal.
@@ 6 Conclusion
We introduce MGSR, a 2D/3D Mutual-boosted Gaussian splatting for SR that enhances both rendering quality and 3D reconstruction accuracy under various light conditions. Moreover, the auto-stop strategy is proposed in the warm-up stage, while geometry-guided illumination decomposition is devised in the mutual-boosted optimization stage. Extensive experiments across various synthetic and real-world datasets at both object and scene levels have validated the superiority of MGSR.
Limitations and future work. Although MGSR achieves reliable reconstruction under reflective conditions, over-smoothness may occur in low-light scenarios, leading to the absence of details. A possible way for addressing this issue is to incorporate exposure compensation for input images, which we will investigate as a future work.
Acknowledgment
This project was partially supported by the Ministry of Education, Singapore, under its Academic Research Fund Grant (RT19/22). The computational experiments were conducted using the CFFF platform at Fudan University.